# Stochastic interpretation of quantum mechanics

### v1

@@ I Introduction
In accordance with the standard interpretation of quantum mechanics omnes1994 ; auletta2001 ; barandes2020 ; freire2022 ; barioni2022 ; pessoa2025 the square of the absolute value of the wave function is related to probability. More precisely, if $\psi(x)$ is the wave function associate to the position $x$ of a particle that moves along a straight line then $|\psi(x)|dx$ is the probability of finding the particle between $x$ and $x+dx$. However, we need also to state the meaning or the interpretation of probability. Within the standard interpretation, it seems that it is interpretated as related to frequency, which can be measured experimentally.
In accordance with the standard interpretation, the probabilistic character of the wave function emerges at the instant of measurement, when the wave function is said to collapse. A question then arises as to the probabilistic character of the wave function prior to measurement. That is, suppose that we perform a measurement at time $t$ and ask whether at a time $t^{\prime}<t$ the wave function has the probabilistic character. To answer this question we must take a measurement at the instant $t^{\prime}$. But, if we do that the wave function collapses at $t$ and the measurement at time $t$ becomes compromised.
From these reasonings, we may say that the probabilistic character of the wave function prior to measurement is not li ble to measurement. Within the standard interpretation of quantum mechanics, it is assumed that the wave function is not associated to probability prior to measurement, and that the probabilistic character emerge at the instant of measurement. Our aim here is to provide an approach in which the wave function does have a probabilistic character at any time. This is acomplished by treating the wave function as a stochastic variable which obeys a stochastic equation of motion oliveira2024a ; oliveira2024b ; oliveira2025 . It is important to say that the probability character at the instant of measurement given by the present approach is the same as that given by the standard interpretation.
@@ VI Continuous space
where $\phi_{j}(x)$ are the eigenfunctions of a certain diferential operator which forms a complete set of orthonormalized functions. The $c_{j}$ are coeffients which depend on time and are understood as stochastic variables.
@@
where here $\hat{H}$ is the Hamiltonian differentical operator and $\psi^{\prime}$ is the wave function at time $t+\tau$. Using the same reasoning as before, we conclude that the norm of the stochastic wave function is preserved at any stochastic trajectory,
@@
Equation (41) becomes identical to (10) if we formally replace $c_{j}$ by $z_{j}$. Therefore we may draw the same conclusions. For instante, the equation (26) becomes
@@
which identify with the Schrödinger equation. Notice that $\psi_{s}(x)$ should no be confused with $\psi(x)$ given by (37) which we call stochastic wave function because the coefficients $c_{j}$ are stochastic variables. The solutions of the type (51) allow us to write $\rho_{jk}=\langle c_{j}c_{k}^{\rangle}=a_{j}a_{k}^{*}$, in which case the Schrödinger wave function is written as

### v2

@@ I Introduction
In accordance with the standard interpretation of quantum mechanics omnes1994 ; auletta2001 ; barandes2020 ; freire2022 ; barioni2022 ; pessoa2025 the square of the absolute value of the wave function is related to probability. More precisely, if $\psi(x)$ is the wave function associate to the position $x$ of a particle that moves along a straight line then $|\psi(x)|dx$ is the probability of finding the particle between $x$ and $x+dx$. However, we need also to state the meaning or the interpretation of probability. Within the standard interpretation, it seems that it is interpreted as related to frequency, which can be measured experimentally.
In accordance with the standard interpretation, the probabilistic character of the wave function emerges at the instant of measurement, when the wave function is said to collapse. A question then arises as to the probabilistic character of the wave function prior to measurement. That is, suppose that we perform a measurement at time $t$ and ask whether at a time $t^{\prime}<t$ the wave function has the probabilistic character. To answer this question we must take a measurement at the instant $t^{\prime}$. But, if we do that the wave function collapses at $t^{\prime}$ and the measurement at time $t$ becomes compromised.
From these reasonings, we may say that the probabilistic character of the wave function prior to measurement is not liable to measurement. Within the standard interpretation of quantum mechanics, it is assumed that the wave function is not associated to probability prior to measurement, and that the probabilistic character emerge at the instant of measurement. Our aim here is to provide an approach in which the wave function does have a probabilistic character at any time. This is accomplished by treating the wave function as a stochastic variable which obeys a stochastic equation of motion oliveira2024a ; oliveira2024b ; oliveira2025 . It is important to say that the probability character at the instant of measurement given by the present approach is the same as that given by the standard interpretation.
Our stochastic approach should not be confused or identified with previous stochastic approaches nelson1966 in which the stochastic variable is the quantum position of the particle and not the wave function. Notice that in our approach the number of stochastic variable is proportional to the dimension of the Hilbert space whereas in these previous stochastic approaches it is equal to the number of degrees of freedom of the quantum system, which is one for a quantum particle moving on a straight line.
From our approach it follows that the covariance matrix associated to the stochastic wave function is identified as the quantum density matrix. The equation that one obtains for the covariances, which is a closed equation in the covariances, is then identified with the quantum Liouville equation. This equation may have solutions consisting of the product of a function and its complex conjugate. When this happens these functions are identified as the pure states.
The stochastic equations that we use are the classical Hamiltonian equations of motion complemented by a noise. The noise we use is a special type of noise that preserves the norm of the stochastic wave function. This is not merely a technical detail but has a conceptual significance because without it we would not obtain quantization properties from classical motion.
We are interested here only in the part of the interpretation of quantum mechanics associated with the probabilistic aspect considering our stochastic treatment of the wave function. This treatment allows to understand that the probabilistic character of quantum mechanics occurs at all times, and not just at the time of measurement, thus making the collapse of the wave function unnecessary.
@@ VI Continuous space
where $\phi_{j}(x)$ are the eigenfunctions of a certain differential operator which forms a complete set of orthonormalized functions. The $c_{j}$ are coefficients which depend on time and are understood as stochastic variables.
@@
where here $\hat{H}$ is the Hamiltonian differential operator and $\psi^{\prime}$ is the wave function at time $t+\tau$. Using the same reasoning as before, we conclude that the norm of the stochastic wave function is preserved at any stochastic trajectory,
@@
Equation (41) becomes identical to (10) if we formally replace $c_{j}$ by $z_{j}$. Therefore we may draw the same conclusions. For instance, the equation (26) becomes
@@
which identify with the Schrödinger equation. Notice that $\psi_{s}(x)$ should no be confused with $\psi(x)$ given by (37) which we call stochastic wave function because the coefficients $c_{j}$ are stochastic variables. The solutions of the type (51) allow us to write $\rho_{jk}=\langle c_{j}c_{k}^{*}\rangle=a_{j}a_{k}^{*}$, in which case the Schrödinger wave function is written as